A storage tank contains 18.5 kg of nitrogen (N_2) at an absolute pressure of 3.00 atm. What will the pressure be if the nitrogen is replaced by an equal mass of CO_2?
Answer #1

From ideal gas equation,

PV = nRT

(P/n) = RT/V

No. of moles of N2 = 18.5 x 1000/28 = 660.71 moles

No. of moles of CO2 = 18.5 x 1000/44 = 420.45 moles

(3/660.71) = P'/420.45

P' = 1.91 atm

Comment in case any doubt please rate my answer.....